OpenStudy (cedric):

An architect needs to determine the slope between two points on a ski lift. The two points have been identified as (25, 30) and (125, 50), where x is the horizontal distance and y is the vertical distance from the bottom of the lift. Assuming the lift runs in a straight line, what is the slope of the line between the two points? (Write your answer in simplest form, using / for a fraction bar if needed.)

hero (hero):

Using y = mx + b, substitute the points to form two equations: 30 = 25m + b 50 = 125m + b Isolate b in each equation: b = 30 - 25m b = 50 - 125m Set b = b, then solve for m: 30 - 25m = 50 - 125m 125m - 25m = 50 - 30 100m = 20 m = 20/100 m = 1/5

hero (hero):

Remember, m means slope in y = mx + b, therefore the slope is 1/5